Find the value of :                                 square root of (-3+4i)
Sorry this sentence is just to incrase characters.

Grade:11

2 Answers

Vikas TU
7 years ago
let 
square root of (-3+4i) = x + iy
squaring both sides,
-3 + 4i = x^2 – y^2 +2xyi
comparing both sides we get,
x^2 – y^2 = – 3................(1)
and
2xy = 4
xy = 2...................(2)
solve both eqns and get x and y values.
Divyansh Thada
7 years ago
@Vikas TU
please write the correct in place of that solution .
again , this line is to incrase the number of characters .